OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1.

OpenStudy (anonymous):

2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (helder_edwin):

P1: 2=1(1+1) Pk: 2+4+6+...+2k=k(k+1) P(k+1): 2+4+6+...+2k+2(k+1)=(k+1)[(k+1)+1]

OpenStudy (anonymous):

thank you very much

OpenStudy (helder_edwin):

u r welcome